13 March 1995

Mr. Harry Stubbs 12 Thompson Lane Milton, MA 02187

Dear Harry,

Hello!, and greetings from the MITSFS, those desirous of �ne sunshine and wonderful Blue Hills!

Apologies for taking so long in getting around to writing you; I spoke to you brie�y at Arisia about picnic planning...

Anyway, as usual we’d like you to pick the date for the Annual Picnic. May the 7th and 14th look good, though the 30th of April could also work.

As this is such an important and momentous occasion, the Society is anxiously awaiting your consideration and decision. We know that we can count on you to select an appropriate day based on the criteria we hold so dear, viz. Weather, Weather, and not-to-forget-the-most-important, Weather.

Yours, respectfully awaiting reply,

John Hawkinson MITSFS Picnicomm